OpenStudy (anonymous):

olve for x: 3x - 5 = 2x + 6

OpenStudy (lgbasallote):

3x - 5 = 2x +6 add 5 to both sides 3x - 5 + 5 = 2x + 6 + 5 3x = 2x + 11 subtract 2x from both sides 3x - 2x = 2x - 2x + 11 2x - 2x = 11

OpenStudy (lgbasallote):

that last line is 3x - 2x = 11

OpenStudy (anonymous):

x=11 answer

hero (hero):

@HillDP, please do not give answers without explanations

OpenStudy (anonymous):

hmms, im sorry. okay @Hero.

OpenStudy (anonymous):

maybe next time. i will. =]]